OpenStudy (anonymous):

Which value is a solution of the equation 5 – 4x = –3 (Medals given to best answer) Which value is a solution of the equation 5 – 4x = –3 0 2 -3 1/4 Which ordered pair is a solution of the equation y = 4x (1, 3) (–1, –4) (–4, –1) (1, –4)

OpenStudy (anonymous):

@sangya21 hey girl can you help with this

OpenStudy (anonymous):

5-4x = -3 5+3 = 4x solve for x Put each coordinate in equation to check which one satisfies the equation

OpenStudy (anonymous):

omg ty you're a life saver bless you

OpenStudy (anonymous):

Anytime.

OpenStudy (triciaal):

@sangya21 you started out fine what if the choices were not given? first isolate the variable to one side of the equation divide by the coefficient of the variable to make it 1 whatever you do to one side of the equation also do to the other to keep the balance here 5 + 3 = 4x 4 x = 8 divide each side by 4 x = 2

OpenStudy (triciaal):

@sangya21 sorry-- I read "to solve for x " not realizing that you just didn't finish and the rest was for the second question.

OpenStudy (anonymous):

x equals 2